# Generalized Neural Sorting Networks with Error-Free Differentiable Swap Functions

### v1

@@ Abstract
Sorting is a fundamental operation of all computer systems, having been a long-standing significant research topic. Beyond the problem formulation of traditional sorting algorithms, we consider sorting problems for more abstract yet expressive inputs, e.g., multi-digit images and image fragments, through a neural sorting network. To learn a mapping from a high-dimensional input to an ordinal variable, the differentiability of sorting networks needs to be guaranteed. In this paper we define a softening error by a differentiable swap function, and develop an error-free swap function that holds non-decreasing and differentiability conditions. Furthermore, a permutation-equivariant Transformer network with multi-head attention is adopted to capture dependency between given inputs and also leverage its model capacity with self-attention. Experiments on diverse sorting benchmarks show that our methods perform better than or comparable to baseline methods.
@@ 1 Introduction
Traditional sorting algorithms (Cormen et al., 2022), e.g., bubble sort, insertion sort, and quick sort, are a well-established approach to arranging given instances in computer science. Since such a sorting algorithm is a basic component to build diverse computer systems, it has been a long-standing significant research area in science and engineering, and the sorting networks (Knuth, 1998; Ajtai et al., 1983), which are structurally designed as an abstract device with a fixed number of wires (i.e., a connection for a single swap operation), have been widely used to perform a sorting algorithm on computing hardware.
Given an unordered sequence of $n$ elements $\mathbf{s}=[s_{1},\ldots,s_{n}]\in\mathbb{R}^{n}$, the problem of sorting is defined as to find a permutation matrix $\mathbf{P}\in\{0,1\}^{n\times n}$ that transforms $\mathbf{s}$ into an ordered sequence $\mathbf{s}_{\mathrm{o}}$:
@@
We generalize the formulation of traditional sorting problems to handle more diverse and expressive types of inputs, e.g., multi-digit images and image fragments, which can contain ordinal information semantically. To this end, we extend the sequence of numbers $\mathbf{s}$ to the sequence of vectors $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $d\gg 1$ is an input dimensionality, and consider the following:
@@
where $\mathbf{X}_{\mathrm{o}}$ and $\mathbf{X}$ are ordered and unordered inputs, respectively. This generalized sorting problem can be reduced to (1) if we are given a proper mapping $g$ from an input $\mathbf{x}\in\mathbb{R}^{d}$ to an ordinal value $s\in\mathbb{R}$. Without such a mapping $g$, predicting $\mathbf{P}$ in (3) remains more challenging than in (1) because $\mathbf{x}$ is often a highly implicative high-dimensional variable. We address this generalized sorting problem by learning a neural sorting network together with mapping $g$ in an end-to-end manner, given training data $\{\mathbf{X}^{(i)},\mathbf{P}_{\mathrm{gt}}^{(i)})\}_{i=1}^{N}$. The main challenge is to make the whole network $f([g(\mathbf{x}_{1}),\ldots,g(\mathbf{x}_{n})])$ with mapping and sorting differentiable to be effectively trained with a gradient-based learning scheme, which is not the case in general. There has been recent research (Grover et al., 2019; Cuturi et al., 2019; Blondel et al., 2020; Petersen et al., 2021; 2022) to tackle the differentiability issue for such a composite function.
In this paper, following a sorting network-based sorting algorithm with differentiable swap functions (DSFs) (Petersen et al., 2021; 2022), we first define a softening error by a sorting network, which indicates a difference between original and smoothed elements. Then, we propose an error-free DSF that resolves an error accumulation problem induced by a soft DSF; this allows us to guarantee a zero error in mapping $\mathbf{X}$ to a proper ordinal value. Based on this, we develop the sorting network with error-free DSFs where we adopt a permutation-equivariant Transformer architecture with multi-head attention (Vaswani et al., 2017) to capture dependency between high-dimensional inputs and also leverage the model capacity of the neural network with a self-attention scheme.
@@
We will make codes publicly available upon publication.
@@ 2 Sorting Networks with Differentiable Swap Functions
where $x^{\prime}=\min(x,y)$ and $y^{\prime}=\max(x,y)$, which makes the order of $x$ and $y$ correct. For example, if $x>y$, then $x^{\prime}=y$ and $y^{\prime}=x$. Without loss of generality, we can express $\min(\cdot,\cdot)$ and $\max(\cdot,\cdot)$ to the following equations:
@@
where $\sigma(\cdot)$ is differentiable. In addition to its differentiability, either (5) or (6) can be achieved with a sigmoid function $\sigma(x)$, i.e., a $s$-shaped function, which satisfies the following properties that (i) $\sigma(x)$ is non-decreasing, (ii) $\sigma(x)=1$ if $x\to\infty$, (iii) $\sigma(x)=0$ if $x\to-\infty$, (iv) $\sigma(0)=0.5$, and (v) $\sigma(x)=1-\sigma(-x)$. Also, as discussed by Petersen et al. (2022), the choice of $\sigma$ affects the performance of neural network-based sorting network in theory as well as in practice. For example, an optimal monotonic sigmoid function, visualized in Figure 4, is defined as
@@
By either (5) or (6), the permutation matrix $\mathbf{P}$ (henceforth, denoted as $\mathbf{P}_{\mathrm{hard}}$ and $\mathbf{P}_{\mathrm{soft}}$ for (5) and (6), respectively) is calculated by the procedure of sorting network:
  1. 1.
    Building a pre-defined sorting network with a fixed number of wires – a wire is a component for comparing and swapping two elements;
  2. 2.
    Feeding an unordered sequence $\mathbf{s}$ into the pre-defined sorting network and calculating a wire-wise permutation matrix $\mathbf{P}_{i}$ for each wire $i$ iteratively;
  3. 3.
    Calculating the permutation matrix $\mathbf{P}$ by multiplying all wire-wise permutation matrices.
As shown in Figure 5, a set of wires is operated simultaneously, so that each set produces an intermediate permutation matrix $\mathbf{P}_{i}$ at $i$th step. Consequently, $\mathbf{P}^{\top}=\mathbf{P}_{1}^{\top}\mathbf{P}_{2}^{\top}\cdots\mathbf{P}_{k}^{\top}=(\mathbf{P}_{k}\cdots\mathbf{P}_{2}\mathbf{P}_{1})^{\top}$ where $k$ is the number of wire sets, e.g., in Figure 5, $k=5$.
@@
Figure 1: Comparisons of diverse DSFs where a swap function is applied once. After a single operation, two input values $x$ and $y$ are softened while our error-free DSF does not change two values. If a distance between $x$ and $y$ is small, softening is more significant.
Figure 1: Comparisons of diverse DSFs where a swap function is applied once. After a single operation, two input values $x$ and $y$ are softened while our error-free DSF does not change two values. If a distance between $x$ and $y$ is small, softening is more significant.
Figure 2: $\mathrm{acc}_{\mathrm{em}}$ (solid) and $\mathrm{acc}_{\mathrm{ew}}$ (dashed), versus sequence lengths. Without a mapping $g$ from $\mathbf{x}$ to $s$, we uniformly sample $n$ elements from $[-10,10]$, which are considered as ordinal scores $\mathbf{s}$. Then, we sort them using the respective sorting networks.
@@ 3 Error-Free Differentiable Swap Functions
Due to the nature of $\overline{\mathrm{min}}$ and $\overline{\mathrm{max}}$, described in (6), the monotonic DSF changes original input values. For example, if $x<y$, then $x<\overline{\mathrm{min}}(x,y)$ and $\overline{\mathrm{max}}(x,y)<y$ after applying the swap function. It can be a serious problem because a change by the DSF is accumulated as the DSF applies iteratively, called an error accumulation problem in this paper. The results of sigmoid functions such as logistic, logistic with ART, reciprocal, Cauchy, and optimal monotonic functions, and also our error-free DSF are presented in Figure 2, where a swap function is applied once; see the work (Petersen et al., 2022) for the respective sigmoid functions. All DSFs except for our error-free DSF change two values, so that they make two values not distinguishable. In particular, if a difference between two values is small, the consequence of softening is more significant than a case with a large difference. Moreover, if we apply a swap function repeatedly, they eventually become identical; see Figure 6 in Section C. While a swap function is not applied as many as it is tested in the synthetic example, it can still cause the error accumulation problem with a few operations. Here we formally define a softening error, which has been mentioned in this paragraph:
@@
Therefore, we define a difference between the original and the softened values, $x^{\prime}-x$ or $y-y^{\prime}$:
@@
As mentioned in the proof of Proposition 2 and empirically shown in Figure 2, a swap function with relatively large $\nabla_{x}\sigma(x)$ changes the original values $x,y$ significantly compared to a swap function with relatively small $\nabla_{x}\sigma(x)$ – they tend to become identical with a small number of operations in the case of large $\nabla_{x}\sigma(x)$.
In addition to the error accumulation problem, such a DSF depends on the scale of $x-y$ as shown in Figure 2. If $x<y$ but $x$ and $y$ are close enough, $\sigma(y-x)$ is between $0.5$ to $1$, which implies that the error can be induced by the scale of $x-y$ (or $y-x$) as well.
@@
Note that $\mathrm{sg}$ indicates that gradients are stopped amid backward propagation, inspired by a straight-through estimator (Bengio et al., 2013). At a step for forward propagation, the error-free DSF produces $x^{\prime}=\min(x,y)$ and $y^{\prime}=\max(x,y)$. On the contrary, at a step for backward propagation, the gradients of $\overline{\mathrm{min}}$ and $\overline{\mathrm{max}}$ are used to update learnable parameters. Consequently, our error-free DSF does not smooth the original elements as shown in Figure 2 and ours shows the 100% accuracy for $\mathrm{acc}_{\mathrm{em}}$ and $\mathrm{acc}_{\mathrm{ew}}$ (see Section 5 for their definitions) as shown in Figure 2. Compared to our DSF, existing DSFs do not correspond the original elements to the elements that have been compared and fail to achieve reasonable performance as a sequence length increases, in the cases of Figure 2.
@@
$x^{\prime}$ $=\left((x\lfloor\sigma_{yx}\rceil+y\lfloor\sigma_{xy}\rceil)-(x\sigma_{yx}+y\sigma_{xy})\right)_{\mathrm{sg}}+(x\sigma_{yx}+y\sigma_{xy})$
$=x\left((\lfloor\sigma_{yx}\rceil-\sigma_{yx})_{\mathrm{sg}}+\sigma_{yx}\right)+y\left((\lfloor\sigma_{xy}\rceil-\sigma_{xy})_{\mathrm{sg}}+\sigma_{xy}\right),$ (12)
$y^{\prime}$ $=x\left((\lfloor\sigma_{xy}\rceil-\sigma_{xy})_{\mathrm{sg}}+\sigma_{xy}\right)+y\left((\lfloor\sigma_{yx}\rceil-\sigma_{yx})_{\mathrm{sg}}+\sigma_{yx}\right),$ (13)
which can be used to define a permutation matrix with the error-free DSF, where $\sigma_{xy}$ and $\sigma_{yx}$ are $\sigma(x-y)$ and $\sigma(y-x)$, respectively. For example, if $n=2$, a permutation matrix $\mathbf{P}$ over $[x,y]$ is
$\mathbf{P}=\left[\begin{matrix}(\lfloor\sigma_{yx}\rceil-\sigma_{yx})_{\mathrm{sg}}+\sigma_{yx}&(\lfloor\sigma_{xy}\rceil-\sigma_{xy})_{\mathrm{sg}}+\sigma_{xy}\\ (\lfloor\sigma_{xy}\rceil-\sigma_{xy})_{\mathrm{sg}}+\sigma_{xy}&(\lfloor\sigma_{yx}\rceil-\sigma_{yx})_{\mathrm{sg}}+\sigma_{yx}\\ \end{matrix}\right].$ (14)
@@
By (11), the softening error $x^{\prime}-\min(x,y)$ (or $\max(x,y)-y^{\prime}$) for an error-free DSF is zero.
@@
Figure 3: Illustration of our neural sorting network with error-free DSFs. Given high-dimensional inputs $\mathbf{X}$, a permutation-equivariant network produces a vector of ordinal variables $\mathbf{s}$, which is used to be swapped using soft or hard sorting network.
@@ 4 Neural Sorting Networks with Error-Free Differentiable Swap Functions
In this section we build a generalized neural network-based sorting network with the error-free DSF and a permutation-equivariant neural network, considering the properties covered in Section 3.
@@
First, we describe a procedure for transforming a high-dimensional input to an ordinal score. Such a mapping $g:\mathbb{R}^{d}\to\mathbb{R}$, which consists of a set of learnable parameters, has to satisfy a permutation-equivariant property:
$[g(\mathbf{x}_{\pi_{1}}),\ldots,g(\mathbf{x}_{\pi_{n}})]=\pi([g(\mathbf{x}_{1}),\ldots,g(\mathbf{x}_{n})]),$ (15)
@@
where $\pi_{i}=[\pi([1,\ldots,n])]_{i}\ \forall i\in[n]$, for any permutation function $\pi$. Typically, an instance-wise neural network, which is applied to each element in a sequence given, is permutation-equivariant (Zaheer et al., 2017). Based on this consequence, instance-wise CNNs are employed in differentiable sorting algorithms (Grover et al., 2019; Cuturi et al., 2019; Petersen et al., 2021; 2022). However, such an instance-wise architecture is limited since it is ineffective for capturing essential features from a sequence. Some types of neural networks such as long short-term memory (Hochreiter & Schmidhuber, 1997) and the standard Transformer architecture (Vaswani et al., 2017) are capable of modeling a sequence of instances, utilizing recurrent connections, scaled dot-product attention, or parameter sharing across elements. While they are powerful for modeling a sequence, they are not obviously permutation-equivariant. Unlike such permutation-variant models, we propose a robust Transformer-based network that satisfies the permutation-equivariant property, which is inspired by the recent work (Vaswani et al., 2017; Lee et al., 2019).
@@
To present our network, we briefly introduce a scaled dot-product attention and multi-head attention:
@@
$\mathrm{att}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{m}}}\right)\mathbf{V}\quad\textrm{and}\quad\mathrm{mha}(\mathbf{Q},\mathbf{K},\mathbf{V})=[\mathrm{head}_{1},\ldots,\mathrm{head}_{h}]\mathbf{W}_{o},$ (16)
@@
where $\mathrm{head}_{i}=\mathrm{att}(\mathbf{Q}\mathbf{W}_{q}^{(i)},\mathbf{K}\mathbf{W}_{k}^{(i)},\mathbf{V}\mathbf{W}_{v}^{(i)})$, $\mathbf{Q},\mathbf{K},\mathbf{V}\!\in\!\mathbb{R}^{n\times hd_{m}}$, $\mathbf{W}_{q}^{(i)},\mathbf{W}_{k}^{(i)},\mathbf{W}_{v}^{(i)}\!\in\!\mathbb{R}^{hd_{m}\times d_{m}}$, and $\mathbf{W}_{o}\!\in\!\mathbb{R}^{hd_{m}\times hd_{m}}$. Similar to the Transformer network, a series of $\mathrm{mha}$ blocks is stacked with layer normalization (Ba et al., 2016) and residual connections (He et al., 2016), and in this paper $\mathbf{X}$ is processed by $\mathrm{mha}(\mathbf{Z},\mathbf{Z},\mathbf{Z})$ where $\mathbf{Z}=g^{\prime}(\mathbf{X})$ or $\mathbf{Z}$ is the output of previous layer; see Section H for the details of the architectures. Note that $g^{\prime}(\cdot)$ is an instance-wise embedding layer, e.g., a simple fully-connected network or a simple CNN. Importantly, compared to the standard Transformer model, our network does not include a positional embedding, in order to satisfy the permutation-equivariant property; $\mathrm{mha}(\mathbf{Z},\mathbf{Z},\mathbf{Z})$ satisfies (15) for the permutation of $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$ where $\mathbf{Z}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{n}]^{\top}$. The output of our network is $\mathbf{s}$, followed by the last instance-wise fully-connected layer. Finally, as shown in Figure 3, our sorting network is able to produce differentiable permutation matrices over $\mathbf{s}$, i.e., $\mathbf{P}_{\mathrm{hard}}$ and $\mathbf{P}_{\mathrm{soft}}$, by utilizing (11) and (6), respectively. Note that $\mathbf{P}_{\mathrm{hard}}$ and $\mathbf{P}_{\mathrm{soft}}$ are doubly-stochastic by Proposition 1.

### v2

@@ Abstract
Sorting is a fundamental operation of all computer systems, having been a long-standing significant research topic. Beyond the problem formulation of traditional sorting algorithms, we consider sorting problems for more abstract yet expressive inputs, e.g., multi-digit images and image fragments, through a neural sorting network. To learn a mapping from a high-dimensional input to an ordinal variable, the differentiability of sorting networks needs to be guaranteed. In this paper we define a softening error by a differentiable swap function, and develop an error-free swap function that holds a non-decreasing condition and differentiability. Furthermore, a permutation-equivariant Transformer network with multi-head attention is adopted to capture dependency between given inputs and also leverage its model capacity with self-attention. Experiments on diverse sorting benchmarks show that our methods perform better than or comparable to baseline methods.
@@ 1 Introduction
Traditional sorting algorithms (Cormen et al., 2022), e.g., bubble sort, insertion sort, and quick sort, are a well-established approach to arranging given instances in computer science. Since such a sorting algorithm is a basic component to build diverse computer systems, it has been a long-standing significant research area in science and engineering. Moreover, sorting networks (Knuth, 1998; Ajtai et al., 1983), which are structurally designed as an abstract device with a fixed number of wires, have been widely used to perform a sorting algorithm on computing hardware, where each wire corresponds to a connection for a single swap operation.
Given an unordered sequence of $n$ elements $\mathbf{s}=[s_{1},\ldots,s_{n}]\in\mathbb{R}^{n}$, the problem of sorting is defined to find a permutation matrix $\mathbf{P}\in\{0,1\}^{n\times n}$ that transforms $\mathbf{s}$ into an ordered sequence $\mathbf{s}_{\mathrm{o}}$:
@@
We generalize the formulation of traditional sorting problems to handle more diverse and expressive types of inputs, e.g., multi-digit images and image fragments, which can contain ordinal information semantically. To this end, we extend the sequence of scalars $\mathbf{s}$ to the sequence of vectors $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times d}$, where $d\gg 1$ is an input dimensionality, and consider the following:
@@
where $\mathbf{X}_{\mathrm{o}}$ and $\mathbf{X}$ are ordered and unordered inputs, respectively. This generalized sorting problem can be reduced to (1) if we are given a proper mapping $g$ from an input $\mathbf{x}\in\mathbb{R}^{d}$ to an ordinal value $s\in\mathbb{R}$. Without such a mapping $g$, predicting $\mathbf{P}$ in (3) remains more challenging than in (1) because $\mathbf{x}$ is often a highly implicative high-dimensional input. We address this generalized sorting problem by learning a neural sorting network together with a mapping $g$ in an end-to-end manner, given training data $\{(\mathbf{X}^{(i)},\mathbf{P}_{\mathrm{gt}}^{(i)})\}_{i=1}^{N}$. The main challenge is to make the whole network $f([g(\mathbf{x}_{1}),\ldots,g(\mathbf{x}_{n})])$ with mapping and sorting components differentiable in order to effectively train the network with a gradient-based learning scheme, which is not the case in general. To tackle the differentiability issue for such a composite function, there has been recent research (Grover et al., 2019; Cuturi et al., 2019; Blondel et al., 2020; Petersen et al., 2021; 2022).
In this paper, following a sorting network-based sorting algorithm with differentiable swap functions (DSFs) (Petersen et al., 2021; 2022), we first define a softening error by a sorting network, which indicates a difference between original and smoothed elements. Then, we propose an error-free DSF that resolves an error accumulation problem induced by a soft DSF; this allows us to guarantee a zero error in mapping $\mathbf{X}$ to proper ordinal values. Based on this, we develop the sorting network with error-free DSFs where we adopt a permutation-equivariant Transformer architecture with multi-head attention (Vaswani et al., 2017) to capture dependency between high-dimensional inputs and also leverage the model capacity of the neural network with a self-attention scheme.
@@ 2 Sorting Networks with Differentiable Swap Functions
where $x^{\prime}=\min(x,y)$ and $y^{\prime}=\max(x,y)$, which makes the order of $x$ and $y$ correct. For example, if $x>y$, then $x^{\prime}=y$ and $y^{\prime}=x$. Without loss of generality, we can express $\min(\cdot,\cdot)$ and $\max(\cdot,\cdot)$ with the following equations:
@@
where $\sigma(\cdot)$ is differentiable. In addition to its differentiability, either (5) or (6) can be achieved with a sigmoid function $\sigma(x)$, i.e., a $s$-shaped function, which satisfies the following properties that (i) $\sigma(x)$ is non-decreasing, (ii) $\sigma(x)=1$ if $x\to\infty$, (iii) $\sigma(x)=0$ if $x\to-\infty$, (iv) $\sigma(0)=0.5$, and (v) $\sigma(x)=1-\sigma(-x)$. Also, as discussed by Petersen et al. (2022), the choice of $\sigma$ affects the performance of neural network-based sorting network in theory as well as in practice. For example, an optimal monotonic sigmoid function, which is visualized in Figure 4, is defined as the following:
@@
By either (5) or (6), the permutation matrix $\mathbf{P}$ (henceforth, denoted as $\mathbf{P}_{\mathrm{hard}}$ and $\mathbf{P}_{\mathrm{soft}}$ for (5) and (6), respectively) is calculated by the following procedure of a sorting network: (i) Building a pre-defined sorting network with a fixed number of wires – a wire is a component for comparing and swapping two elements; (ii) Feeding an unordered sequence $\mathbf{s}$ into the pre-defined sorting network and calculating a wire-wise permutation matrix $\mathbf{P}_{i}$ for each wire $i$ iteratively; (iii) Calculating the permutation matrix $\mathbf{P}$ by multiplying all wire-wise permutation matrices.
As shown in Figure 2, a set of wires represents a set of swap operations that are operated simultaneously, so that each set produces an intermediate permutation matrix $\mathbf{P}_{i}$ at the $i$th step. Consequently, $\mathbf{P}^{\top}=\mathbf{P}_{1}^{\top}\mathbf{P}_{2}^{\top}\cdots\mathbf{P}_{k}^{\top}=(\mathbf{P}_{k}\cdots\mathbf{P}_{2}\mathbf{P}_{1})^{\top}$, where $k$ is the number of wire sets. For example, in Figure 2, $k=5$.
@@
In Sections 3 and 4, we present an error-free DSF and a neural sorting network with error-free DSFs.
Figure 1: A sorting network with 5 wire sets and their permutation matrices.
Figure 1: A sorting network with 5 wire sets and their permutation matrices.
Figure 2: Comparisons of diverse DSFs where a swap function is applied once. After a single operation, two input values $x$ and $y$ are softened while our error-free DSF does not change two values. If $|x-y|$ is small, softening will be more significant.
@@ 3 Error-Free Differentiable Swap Functions
Due to the nature of $\overline{\mathrm{min}}$ and $\overline{\mathrm{max}}$, which is described in (6), the monotonic DSF changes original input values. For example, if $x<y$, then $x<\overline{\mathrm{min}}(x,y)$ and $\overline{\mathrm{max}}(x,y)<y$ after applying the swap function. It can be a serious problem because changes by the DSF are accumulated as the DSF applies iteratively, called an error accumulation problem in this paper. The results of sigmoid functions such as the logistic, logistic with ART, reciprocal, Cauchy, and optimal monotonic functions, and also our error-free DSF are presented in Figure 2, where a swap function is applied once; see the work (Petersen et al., 2022) for the respective sigmoid functions. All DSFs except for our error-free DSF change two values, so that they can make two values not distinguishable. In particular, if a difference between two values is small, the consequence of softening is more significant than a case with a large difference. Moreover, if we apply a swap function repeatedly, they eventually become identical; see Figure 5 in Section B. While a swap function is not applied as many as it is tested in the synthetic example shown in Figure 5, it can still cause the error accumulation problem with a few operations. Here we formally define a softening error, which has been mentioned in this paragraph:
@@
Therefore, we define a difference between the original and softened values, $x^{\prime}-x$ or $y-y^{\prime}$:
@@
As mentioned in the proof of Proposition 2 and empirically shown in Figure 2, a swap function with relatively large $\nabla_{x}\sigma(x)$ changes the original values $x,y$ significantly compared to a swap function with relatively small $\nabla_{x}\sigma(x)$ – they tend to become identical with the small number of operations in the case of large $\nabla_{x}\sigma(x)$.
In addition to the error accumulation problem, such a DSF depends on the scale of $|x-y|$ as shown in Figure 2. If $x<y$ but $x$ and $y$ are close enough, $\sigma(y-x)$ is between $0.5$ and $1$, which implies that the error can be induced by the scale of $|x-y|$ as well.
@@
Note that $\mathrm{sg}$ indicates that gradients are stopped amid backward propagation, inspired by a straight-through estimator (Bengio et al., 2013). At a step for forward propagation, the error-free DSF produces $x^{\prime}=\min(x,y)$ and $y^{\prime}=\max(x,y)$. On the contrary, at a step for backward propagation, the gradients of $\overline{\mathrm{min}}$ and $\overline{\mathrm{max}}$ are used to update learnable parameters. Consequently, our error-free DSF does not smooth the original elements as shown in Figure 2 and our DSF shows 100% accuracy for $\mathrm{acc}_{\mathrm{em}}$ and $\mathrm{acc}_{\mathrm{ew}}$ (see Section 5 for their definitions) as shown in Figure 6. Compared to our DSF, the existing DSFs do not correspond the original elements to the elements that have been compared and fail to achieve reasonable performance as a sequence length increases, in the cases of Figure 6.
@@
$x^{\prime}$ $\!=\!\left((x\lfloor\sigma(y-x)\rceil\!+\!y\lfloor\sigma(x-y)\rceil)\!-\!(x\sigma(y-x)\!+\!y\sigma(x-y))\right)_{\mathrm{sg}}\!+\!(x\sigma(y-x)\!+\!y\sigma(x-y))$
$\!=\!x\left((\lfloor\sigma(y-x)\rceil\!-\!\sigma(y-x))_{\mathrm{sg}}\!+\!\sigma(y-x)\right)\!+\!y\left((\lfloor\sigma(x-y)\rceil\!-\!\sigma(x-y))_{\mathrm{sg}}\!+\!\sigma(x-y)\right),$ (12)
$y^{\prime}$ $\!=\!x\left((\lfloor\sigma(x-y)\rceil\!-\!\sigma(x-y))_{\mathrm{sg}}\!+\!\sigma(x-y)\right)\!+\!y\left((\lfloor\sigma(y-x)\rceil\!-\!\sigma(y-x))_{\mathrm{sg}}\!+\!\sigma(y-x)\right),$ (13)
which can be used to define a permutation matrix with the error-free DSF. For example, if $n=2$, a permutation matrix $\mathbf{P}$ over $[x,y]$ is
$\mathbf{P}=\left[\begin{matrix}(\lfloor\sigma(y-x)\rceil-\sigma(y-x))_{\mathrm{sg}}+\sigma(y-x)&(\lfloor\sigma(x-y)\rceil-\sigma(x-y))_{\mathrm{sg}}+\sigma(x-y)\\ (\lfloor\sigma(x-y)\rceil-\sigma(x-y))_{\mathrm{sg}}+\sigma(x-y)&(\lfloor\sigma(y-x)\rceil-\sigma(y-x))_{\mathrm{sg}}+\sigma(y-x)\\ \end{matrix}\right].$ (14)
@@
By (11), the softening error $x^{\prime}-\min(x,y)$ or $\max(x,y)-y^{\prime}$ for an error-free DSF is zero.
@@
Figure 3: Illustration of our neural sorting network with error-free DSFs. Given high-dimensional inputs $\mathbf{X}$, a permutation-equivariant network produces a vector of ordinal variables $\mathbf{s}$, which is used to be swapped using a soft or hard sorting network.
@@ 4 Neural Sorting Networks with Error-Free Differentiable Swap Functions
In this section we propose a generalized neural network-based sorting network with an error-free DSF and a permutation-equivariant neural network, considering the properties covered in Section 3.
@@
First, we describe a procedure for transforming high-dimensional inputs to ordinal scores. Such a mapping $g:\mathbb{R}^{d}\to\mathbb{R}$, which consists of a set of learnable parameters, has to satisfy a permutation-equivariant property:
$[g(\mathbf{x}_{\pi_{1}}),g(\mathbf{x}_{\pi_{2}}),\ldots,g(\mathbf{x}_{\pi_{n}})]=\pi([g(\mathbf{x}_{1}),g(\mathbf{x}_{2}),\ldots,g(\mathbf{x}_{n})]),$ (15)
@@
where $\pi_{i}=[\pi([1,2,\ldots,n])]_{i}\ \forall i\in[n]$, for any permutation function $\pi$. Typically, an instance-wise neural network, which is applied to each element in a sequence given, is permutation-equivariant (Zaheer et al., 2017). Based on this property, instance-wise CNNs are employed in differentiable sorting algorithms (Grover et al., 2019; Cuturi et al., 2019; Petersen et al., 2021; 2022). However, such an instance-wise architecture is limited since it is ineffective for capturing essential features from a sequence. Some types of neural networks such as long short-term memory (Hochreiter & Schmidhuber, 1997) and the standard Transformer architecture (Vaswani et al., 2017) are capable of modeling a sequence of instances, utilizing recurrent connections, scaled dot-product attention, or parameter sharing across elements. While they are powerful for modeling a sequence, they are not obviously permutation-equivariant. Instead of such permutation-sensitive models, we adopt a robust Transformer-based network that satisfies the permutation-equivariant property, which is inspired by the recent work (Vaswani et al., 2017; Lee et al., 2019).
@@
To explain our network, we briefly introduce scaled dot-product attention and multi-head attention:
@@
$\mathrm{att}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{m}}}\right)\mathbf{V}\quad\textrm{and}\quad\mathrm{mha}(\mathbf{Q},\mathbf{K},\mathbf{V})=[\mathrm{head}_{1},\mathrm{head}_{2},\ldots,\mathrm{head}_{h}]\mathbf{W}_{o},$ (16)
@@
where $\mathrm{head}_{i}=\mathrm{att}(\mathbf{Q}\mathbf{W}_{q}^{(i)},\mathbf{K}\mathbf{W}_{k}^{(i)},\mathbf{V}\mathbf{W}_{v}^{(i)})$, $\mathbf{Q},\mathbf{K},\mathbf{V}\!\in\!\mathbb{R}^{n\times hd_{m}}$, $\mathbf{W}_{q}^{(i)},\mathbf{W}_{k}^{(i)},\mathbf{W}_{v}^{(i)}\!\in\!\mathbb{R}^{hd_{m}\times d_{m}}$, and $\mathbf{W}_{o}\!\in\!\mathbb{R}^{hd_{m}\times hd_{m}}$. Similar to the Transformer network, a series of $\mathrm{mha}$ blocks is stacked with layer normalization (Ba et al., 2016) and residual connections (He et al., 2016), and in this paper $\mathbf{X}$ is processed by $\mathrm{mha}(\mathbf{Z},\mathbf{Z},\mathbf{Z})$ where $\mathbf{Z}=g^{\prime}(\mathbf{X})$ or $\mathbf{Z}$ is the output of a previous layer; see Section I for the details of the architectures. Note that $g^{\prime}(\cdot)$ is an instance-wise embedding layer, e.g., a simple fully-connected network or a simple CNN. Importantly, compared to the standard Transformer model, our network does not include a positional embedding, in order to satisfy the permutation-equivariant property; $\mathrm{mha}(\mathbf{Z},\mathbf{Z},\mathbf{Z})$ satisfies (15) for the permutation of $\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{n}$ where $\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{n}]^{\top}$. The output of our network is $\mathbf{s}$, followed by the last instance-wise fully-connected layer. Finally, as shown in Figure 3, our sorting network is able to produce differentiable permutation matrices over $\mathbf{s}$, i.e., $\mathbf{P}_{\mathrm{hard}}$ and $\mathbf{P}_{\mathrm{soft}}$, by utilizing (11) and (6), respectively. Note that $\mathbf{P}_{\mathrm{hard}}$ and $\mathbf{P}_{\mathrm{soft}}$ are doubly-stochastic by Proposition 1. In addition, the details of the permutation-equivariant network with multi-head attention are briefly visualized in Figure 7.